OpenStudy (tgstudios):

simplify 5 square root of 7 + 12 times square root of 6 minus 10 times square root of 7 - 5 times square root of 6.

OpenStudy (tgstudios):

Please I give medal:0

OpenStudy (tgstudios):

\[5\sqrt{7} + 12 \times \sqrt{6} - 10 \times \sqrt{7} - 5 \times \sqrt{6}\] Theres the problem

OpenStudy (tgstudios):

simplify it help

OpenStudy (rcp031):

\[5\sqrt{7}-10\sqrt{7}+12\sqrt{6}-5\sqrt{6} \]

OpenStudy (tgstudios):

ohhh.... ok

OpenStudy (tgstudios):

forgot that... :/

OpenStudy (tgstudios):

help me simplify this but I think I might have it, plz check me

OpenStudy (tgstudios):

I have half of it...

OpenStudy (tgstudios):

I'm so confused I'm sorry I'm being so dumb

OpenStudy (karim728):